Ruthenium-106 (106Ru) is used to treat melanoma in the eye. This isotope decays by β− emission with a half-life of 373.59 days. One source of the isotope is reprocessed nuclear reactor fuel.

How many protons and how many neutrons does theRu106 nucleus contain?

protons:

neutrons:

Identify the missing daughter nucleus in the β− emission decay of Ru106

Ru106⟶ + e−+ν¯e+ e−+ν¯e

How many years after Ru106 is implanted in the eye does it take for its decay rate (or activity) to be reduced to 21.5% of the initial decay rate? Use 1 year=365 daya

time: answer in years
Answer #1

Half life time, ty, = 373.59 days of Ru-los a atomic number (2)= hoof Bocoton (p) Protons 44 Neutrons = 106-44 = 62 ③ Reactio

828.65 days X Given activity A A 100 828-65 days after stimet A=21.51. of to 22:1 Activity equation with hoff dage time Ao et